OpenStudy (anonymous):

Question) Show that Torricelli's Law may be solved by separation of variables. State Assumptions

OpenStudy (anonymous):

can you post Toricelli's law

OpenStudy (anonymous):

\[A(h)\frac{ dh }{ dt } = -k \sqrt{h}\]

OpenStudy (anonymous):

not sure what A(h) stands for

OpenStudy (anonymous):

Toricelli’s Law. Suppose that a water tank has a hole with area \( a\) at its bottom and that water is draining from the hole. Let y(t) (in feet) and V(t ) (in cubic feet) denote the depth and the volume of water in the tank at time t (in seconds). Then (under ideal conditions) the velocity of the stream of water exiting the tank will be \( v = \sqrt{ 2g y }\)

OpenStudy (anonymous):

yeah, I get that part. And I know how velocity in this case is also derived. I'm just curious on the conditions required to make this 'equation' work.

OpenStudy (irishboy123):

\[-\frac{A(h)}{\sqrt{h}} \, dh = -k \, dt \] i posted something on this yesterday, will try find a link, it goes into more detail

OpenStudy (anonymous):

here is a pdf on the separation of variable, and an intro to it

OpenStudy (irishboy123):

down the bottom of this http://openstudy.com/users/irishboy123#/updates/55fffe4ae4b0ed58e276cd99

OpenStudy (anonymous):

Aright, Much appreciated

OpenStudy (anonymous):

@IrishBoy123 can you explain this part in the pdf i posted http://prntscr.com/8j6f7s It follows that V =∫ A(y)*dy <---- this is my reasoning then dV/dt = d/dt ( ∫ A(y)*dy ) = A(t) ?

OpenStudy (anonymous):

oh i think it follows from chain rule

OpenStudy (anonymous):

V =∫ A(y)*dy on [ 0, k] where k is the height of the vessel. dV/dt = d/dt ( ∫ A(y)*dy ) = d/dy ( ∫ A(y)*dy ) * dy/dt = A(y) * dy/dt I am not sure exactly what rule this is. AL

OpenStudy (irishboy123):

yeah, chain rule that pdf is a good find :-)

OpenStudy (anonymous):

i hate it when they skip steps,. this does not seem obvious to me, lol

OpenStudy (anonymous):

I noticed your post had a more interesting argument using volume and washers . but this is pretty cool too . i think i made a proof

OpenStudy (anonymous):

Want to make sure the limits make sense here. Let \( A(y) \) be the cross sectional area at height \( y \) of vessel. $$ { \large V(t) = \int_{0}^{y(t)}A(y)~ dy \\ \frac{dV}{dt}= \frac{d}{dt} \left( \int_{0}^{y(t)}A(y) dy \right)= \frac{d}{dy} \left( \int_{0}^{y(t)}A(y) dy \right)\cdot \frac{dy}{dt}= A(y) \cdot \frac{dy}{dt} } $$

OpenStudy (anonymous):

wikipedia has a nice argument https://en.wikipedia.org/wiki/Torricelli's_law#Application_for_time_to_empty_the_container